OpenStudy (anonymous):

find the rate of change for the situation you run 7 miles in one hour and 14 miles in two hours

OpenStudy (anonymous):

Are you looking for rate of change in speed between the two runs? Or just how fast you run in general?

OpenStudy (paki):

(14-7) / (2-1) = 7/1 = 7mph

OpenStudy (paki):

@somayapona

OpenStudy (anonymous):

there is no change in speed